OpenStudy (anonymous):

for the following simultaneous equations, k is a real constant: kx-3y=k+3 4x+(k+7)y=1 Find the value of k for which there are infinitely many solutions.

Parth (parthkohli):

\[\frac{k}{4} = \frac{-3}{k+7} = \frac{k+3}{1}\]

OpenStudy (anonymous):

i dont understand this

OpenStudy (michele_laino):

Please, as @ParthKohli has suggested, you have to pick the common root of these three equations: \[\begin{gathered} \frac{k}{4} = \frac{{ - 3}}{{k + 7}} \hfill \\ \frac{k}{4} = k + 3 \hfill \\ \frac{{ - 3}}{{k + 7}} = k + 3 \hfill \\ \end{gathered} \]

OpenStudy (thomas5267):

\[ \begin{vmatrix} k&-3\\ 4x&k+7 \end{vmatrix} =0 \] The determinant must be zero as the equations have infinitely many solution. Not absolutely sure though. Check your notes.